OpenStudy (anonymous):

{x-y=5 {2x+y=1 solve by using a matrix equation

jhonyy9 (jhonyy9):

x-y=5 2x+y=1 ---------- 3x =6 x=6/3 x=2

OpenStudy (turingtest):

\[\left[\begin{matrix}1 & -1 \\ 2 & 1\end{matrix}\right]\overrightarrow{x}=\left(\begin{matrix}5 \\ 1\end{matrix}\right)\]solve with Gaussian elimination, Cramer's rule, or whatever you wish.